OpenStudy (anonymous):

When simplifying an expression, you ____ perform operations inside grouping symbols first. A. always B. sometimes C. never

OpenStudy (amistre64):

sometimes

OpenStudy (amistre64):

im sure the answer they want you to give is always; but thats not the case ....

OpenStudy (amistre64):

\[\frac{2(3-1)}{4}=\frac{3-1}{2}\] just as in that example I did not start with the (..) stuff